How to calculate the oxidation number of N in HCN ?

Grade:11

2 Answers

Abhishek Sharma
5 years ago
ON. of H=+1 & C=+4 in HCN .
Let ON. Of N=x
1+x+4=0
x+5=0
X=-5 
so the Ox. No. Of N =-5
Simply we can find the ox. No.of any element in the compound or in any molecule simply by putting the Ox. No. Of othe elements .
john carlos
5 years ago
dear student      
HCN is compound so its oxidation number will be    zero                                                 
H  has  +1  and  C  has  +4  so                               
let  x = N                                                                      
x + 1 + 4 = 0                                                                 
x + 5 = 0
x= -5